OpenStudy (anonymous):

PLEASE HELP? ln ( 1/e^4 )

OpenStudy (anonymous):

\[\ln e^{-4}=\] use \[\ln a^b=b \ln a\]

OpenStudy (anonymous):

write that out as: \(\large ln(\frac{1}{e^4})= ln(e^{-4}) \) then use the properties of logs...